A 1.800-g sample of solid phenol (C6H5OH(s)) was burned in a bomb calorimeter, which has a total heat capacity of 11.66 kJ/∘C. The temperature of the calorimeter plus its contents increased from 21.36∘C to 26.37∘C.

A. Write a balanced chemical equation for the reaction that takes place in the bomb calorimeter.

Express you answer as a chemical equation including phases.

B. What is the heat of combustion per gram of phenol?

Express the heat in kilojoules per gram to three significant digits.

C. What is the enthalpy per mole of phenol?

Express the enthalpy in kilojoules per mole to three significant digits.
Answer #1

A.

C6H5OH(s)   + 7O2(g) -----------------> 6CO2(g)   + 3H2O(l)

B. Total heat capacity of 11.66 kJ/∘C

    T   = T2-T1

              = 26.37-21.36 = 5.010C

The heat of combustion    = Total heat capacity * T  

                                          = 11.66KJ/0C*5.010C = 58.4166KJ

the heat of combustion per gram of phenol    = 58.4166KJ/1.8g = 32.5KJ/g

C.

no of moles of C6H5OH    = W/G.M.Wt

                                           = 1.8/94.11   = 0.01913moles

the enthalpy per mole of phenol    = 58.4166/0.01913    = -3054   = -3.05*10^3 KJ/mole